# A Dimca-Greuel type inequality for foliations

### v1

@@ 3. Proof of Theorem A
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$.
Let $C=\{x=0\}$, $\mathcal{B}_{0}=C+C_{0,1}+C_{1,1}$, $\mathcal{B}_{\infty}=C_{2,1}$, then $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices for $\mathcal{F}$. We have
$\mu_{0}(\mathcal{F},\mathcal{B}_{0})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{0,1})+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+29+21-2+1=54,$
$\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=\mu_{0}(\mathcal{F},C_{2,1})=21,$
and
$T_{0}(\mathcal{F},\mathcal{B}_{0})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{0,1})+\tau_{0}(\mathcal{F},C_{1,1})=5+28+21=54.$
Hence
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},\mathcal{B}_{0})-\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})+1=54-21+1=34,$
and
$\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B}_{0})}=\frac{34}{54}<\frac{4}{3}.$
We conclude the paper with the following question:
Question: Is the inequality (24) true if instead of $T_{p}(\mathcal{F},\mathcal{B}_{0})$ we write $\tau_{p}(\mathcal{F},\mathcal{B}_{0})$?

### v2

@@ 5. Applications
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$. Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We have $GSV_{0}(\mathcal{F},C)=5$, $GSV_{0}(\mathcal{F},C_{1,1})=5$, $\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$, hence $\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B})}=\frac{25}{26}<\frac{4}{3}$.
We conclude the paper with a question:
Question 1: Is the inequality (28) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?